OpenStudy (anonymous):

company overvalues its shares by 25 percent. If the quoted price of a share is $300, what is the true and fair price of the share?

OpenStudy (anonymous):

let the original value be x say, then the given question says that x+x*25/100 = 300 that means x+ x/4 =300 5x=1200 x=1200/5=240 :D I hope you understood and its correct

OpenStudy (anonymous):

\[\frac{300}{1.25}=240 \]